### v1

# Thermalization and Many-Body Zeno Effect in monitored Hamiltonian Dynamics

Abstract
Random quantum states are essential for various applications in quantum information science. Prior approaches of generating genuine random states rely on a large bath to thermalize the system, such that a subsequent measurement on the bath post-selects a random state for the system. To reduce the size of the required bath, we propose an alternative approach based on holographic deep thermalization driven by Hamiltonian evolution, combined with mid-circuit measurements. By trading spatial and time resources, our approach achieves genuine randomness with a bath of constant size that is independent of the system size. We quantify randomness with the frame potential and analyze its dynamics throughout the evolution. Given a total evolution time, as we increase the number of mid-circuit measurements, the frame potential initially decreases exponentially with the number of measurements, due to the mechanism of holographic deep thermalization. Past a critical number of mid-circuit measurements, the frame potential rises again, signaling the onset of the quantum Zeno effect. We provide analytical results for the asymptotic behavior of the frame potential, which are in good agreement with the numerical simulations. Our findings offer practical guidance for generating Haar-random ensembles through Hamiltonian evolution and controlled measurement.
I Introduction
The second law of thermodynamics states that macroscopic complex systems thermalizes–their entropy increases towards its maximum value as time evolves. To resolve its contradiction to the unitarity of quantum physics, numerous efforts have been dedicated to describe equilibration and thermalization of isolated quantum system [1, 2, 3, 4, 5], where chaotic quantum dynamics drives the isolated system to an equilibrium such that local observables appear thermal. The key is that a small system interacts with a large bath system and becomes highly entangled with the bath [6], and at the same time the reduced state of the small system appears thermal. More recently, the notion of deep thermalization [7, 8] provides a further interpretation of such a process via considering quantum measurement on the bath.
Deep thermalization concerns the state ensemble generated by quantum measurement on the bath, given that the system and bath are jointly entangled in a pure state. Conditioned on the measurement result, the system is in a random pure state; collecting conditional states under the different measurement outcomes, one therefore obtains an ensemble of pure states for the system. Refs. [7, 8] show that such measurement-induced ensemble becomes Haar random, characterized by high-order moments (hence the name ‘deep thermalization’), when the system-bath unitary dynamics are sufficiently complex. While a large bath compared to the system is essential for deep thermalization [7, 8] and quantum thermalization [1, 2, 3, 4, 5], a small bath with certain dynamical control also plays an important role in statistical physics.
The most famous example is the ‘Maxwell demon’, a small agent that measures the speed of a particle and selects high energy particles to go through to the hotter side of the system. While originally proposed as a conceptual challenge to statistical physics, it also provides a case where a small agent, with the capability of measurement and control, can change the equilibration of a large system. The essential part to resolve the seemingly violation of second law of thermodynamics is that the agent needs to be reset via dissipation to external environment. One can also provide a description of Maxwell’s demon by the quantum collision model [9], where unitary interactions and system reset alternates. Similar to Maxwell demon, stroboscopic thermalization adopts periodically driven and dissipated small ancilla to thermalize the system [10]. While the quantum collision model [9] establishes the general theory model and the corresponding dynamical equations of quantum trajectory, the generation of an ensemble of quantum states are not well understood.
Holographic deep thermalization (HDT) [11] extends the Maxwell’s demon type of bath-system interaction to the state ensemble generation. Via repeated interaction, measurement and reset, a small bath system still enables the generation of Haar random pure states, allowing a reduction of quantum memory in genuine random state generation. In the initial proposal [11], however, each unitary interaction between the system and bath is assumed to be a fixed unitary sampled from Haar ensemble. The initial theory and experimentation in Ref. [11] focuses on quantum circuit dynamics and the Hamiltonian realization of HDT is not well understood. Indeed, in a Hamiltonian dynamics, the effective realization of each Haar-typical unitary requires a long evolution time and therefore potentially creates a challenge in experiments.
Figure 1: Schematic of the quantum circuit for holographic deep thermalization with Hamiltonian dynamics and multiple mid-circuit measurements.
In this work, we extend the study of holographic deep thermalization to Hamiltonian dynamics and uncover a subtle interplay between thermalization and the quantum Zeno effect [12, 13, 14, 15, 16, 17]. By employing an interacting Hamiltonian and performing intermediate measurements on the bath, we aim to generate Haar-random states. Our findings show that increasing the number of intermediate measurements generally enhances the randomness of the resulting state ensemble. However, beyond a certain threshold, excessive measurements trigger the quantum Zeno effect, inhibiting the system’s evolution and diminishing the ensemble’s randomness. We provide both analytical and numerical insights to determine the optimal number of measurements that maximizes Haar-randomness, given a fixed total evolution time.
II Overview
Hamiltonian evolution can also be used to generate Haar-random states and is often more practical to implement in quantum simulation platforms. In our approach, we utilize a Hamiltonian that satisfies the eigenstate-thermalization-hypothesis (ETH) to drive the system toward Haar randomness. The entire system consists of the system with $n_{\rm s}$ qubits and the bath (ancillary) with $n_{\rm b}$ qubits. The total evolution time is fixed at $T$, during which we insert $n$ intermediate measurements on the bath qubits. As a result, each segment of unitary evolution lasts for $\Delta t=T/n$. The bath is measured and then reset after each segment. A schematic of the quantum circuit is shown in Fig. 1. Different measurement outcomes yield different output states, collectively forming an ensemble. We adopt the frame potential to characterize the randomness of the ensemble quantitatively.
For the Hamiltonian evolution process shown in Fig. 1, the frame potential initially decreases and then increases depending on the number of measurements $n$, as illustrated by the blue curves in Fig. 2. The initial decrease in the small-$n$ regime is well captured by the HDT theory, shown as the orange curves in Fig. 2. In contrast, the increase in the large-$n$ regime can be explained by the quantum Zeno effect, represented by the green curves. These theoretical descriptions hold across different bath sizes, as demonstrated by the comparison between cases with a small number of bath qubits in Figs. 2(a)(d) and those with a larger number of bath qubits in Figs. 2(b)(c).
When the total evolution time $T$ is large, the frame-potential decrease saturates once $n$ approaches $n_{\rm sat}\sim O\left(\left(n_{\rm s}-{\rm log}_{2}r\right)/n_{\rm b}\right)$, as indicated by the orange dashed lines in Fig. 2(a)(b), where $r$ is a Hamiltonian dependent parameter. Beyond this point, the frame potential first saturates to a plateau, with fluctuations induced by dynamical revival from finite-system-size effect (see Appendix A). After the plateau, the frame potential begins to increase due to the onset of the quantum Zeno effect, occurring around $n_{\rm Zeno}\sim O\left(\sqrt[\alpha]{T^{\alpha+1}/n_{\rm s}}\right)$, as indicated by the green dashed lines in Fig. 2(a)(b), where $\alpha$ is another Hamiltonian depended parameter.
For small $T$ with $n_{\rm sat}>n_{\rm Zeno}$ (Fig. 2(c)(f)), HDT and the Zeno effect compete, and the minimum frame potential is reached near $n_{\gamma}\sim O\left(T/\sqrt[\alpha+1]{n_{\rm b}}\right)$ which reflects the optimal balance between thermalization and measurement-induced suppression.
III Set-up
In HDT, both the system and bath are initialized in a trivial state that is easy to prepare; here, we choose the all-zero state, denoted as $|\psi_{0}\rangle\otimes|\phi_{0}\rangle$. The system and bath are then entangled through evolution under a time-independent Hamiltonian $H$. They jointly evolve for a time interval $\Delta t$ according to the Schrödinger equation. For brevity, we set $\hbar=1$. After each evolution step, the bath is measured in the computational basis $|\phi_{m}\rangle$} and reset to the trivial state $|\phi_{0}\rangle$. The cycle of Hamiltonian evolution, measurement, and reset is repeated $n$ times, with the total evolution time fixed at $T$. At the end of the process, the system collapses into one of the possible states $|\psi_{z}\rangle$ with probability $p_{z}$, where $z$ indexes all possible measurement trajectories. These outcomes define an ensemble $\mathcal{E}=\{p_{z},|\psi_{z}\rangle\}$.
Figure 2: Frame potential as a function of the number of measurements. The blue lines represent numerical results, while the orange and green lines correspond to analytical results for the asymptotic region of HDT in Eq. (7) and the Zeno effect in Eq. (12), respectively. Other parameters are: (a),(d) 5 system qubits and 3 bath qubits; (b),(c),(e),(f) 7 system qubits and 1 bath qubit. Total evolution time is $T=15$ for (a),(b),(d),(e) and $T=5$ for (c),(f). The orange, green, and blue dashed lines indicate the positions of $n_{\rm sat}$, $n_{\rm Zeno}$, and $n_{\gamma}$, respectively, with $r=0.1$
The randomness of the ensemble $\mathcal{E}$ can be quantified using the frame potential, defined as
$F^{(K)}=\sum_{z,z^{\prime}}p_{z}p_{z^{\prime}}\left|\langle\psi_{z}|\psi_{z^{\prime}}\rangle\right|^{2K}.$ (1)
In the ideal case, the Haar random frame potential is
$F_{\rm Haar}^{(K)}=\frac{(N_{\rm s}-1)!K!}{(N_{\rm s}+K-1)!},$ (2)
where $N_{\rm s}=2^{n_{\rm s}}$ is the size of system Hilbert space.
In prior work on DT, the transverse-field Ising model has been used to approximate Haar-random dynamics. In our approach, we adopt the same Hamiltonian to implement HDT, which is given by
$H_{\rm Ising}=J_{x}\sum\sigma_{j}^{x}+J_{z}\sum\sigma_{j}^{z}+J_{zz}\sum\sigma^{z}_{j}\sigma_{j+1}^{z},$ (3)
where $J_{x}$ denotes the strength of the transverse field, $J_{z}$ the longitudinal field, and $J_{zz}$ the nearest-neighbor interaction strength.
For the more general cases, we introduce strong perturbation terms to the Hamiltonian in order to disrupt any potential underlying structure. Specifically, we consider the following two types of perturbations
$H_{YY}$ $=$ $H_{\rm Ising}+J_{yy}\sum\sigma^{y}_{j}\sigma_{j+1}^{y},$ (4)
$H_{XXX}$ $=$ $H_{\rm Ising}+J_{xxx}\sum\sigma_{j-1}^{x}\sigma^{x}_{j}\sigma_{j+1}^{x}.$ (5)
In the context of Hamiltonian evolution, we consider a system of qubits where a subset of them is treated as a bath and measured, while the remaining qubits form an ensemble that approaches Haar-random. Generally, the Hamiltonian can be written into three parts
$H$ $=H_{{\rm s}}+H_{{\rm b}}+H_{{\rm c}},$ (6)
where $H_{{\rm s}},H_{{\rm b}}$ and $H_{{\rm c}}$ correspond to the system, bath and interaction Hamiltonian.
Although the Hamiltonian may possess symmetries, such as inversion or translational symmetry, which constrain the spreading of the wave function across the full Hilbert space, the measurement and subsequent reset to the bath partially break these symmetries. This symmetry breaking, in turn, enhances both randomness and thermalization.
IV Holographic deep thermalization
Figure 3: The convergence dynamics of frame potential as a function of the number of measurements for (a) $H_{\rm Ising}$, (b) $H_{YY}$, and (c) $H_{XXX}$. y axis is the frame potential difference with the converged value in Eq. (8). Blue lines represent numerical results, while orange lines correspond to analytical predictions for HDT. The total evolution time is $T=15$. The system consists of 7 qubits, with 1 bath qubit.
In the context of holographic deep thermalization, the insertion of multiple mid-circuit measurements significantly accelerates the thermalization process, potentially leading to exponential speedup in approaching Haar randomness. If the unitaries are randomly sampled from the Haar measure, the expected first-order frame potential in holographic deep thermalization is given by
$F^{(1)}=q_{\rm 1}+\frac{(N_{\rm s}-1)(N_{\rm s}N_{\rm b}-1)}{N_{\rm s}^{2}N_{\rm b}+1}\left(\frac{(N_{\rm s}^{2}-1)N_{\rm b}}{N_{\rm s}^{2}N_{\rm b}^{2}-1}\right)^{n},$ (7)
where the corrected saturation value is
$q_{\rm 1}=\frac{N_{\rm s}^{2}(N_{\rm b}+1)}{N_{\rm s}^{2}N_{\rm b}+1}F_{\rm Haar}^{(1)}.$ (8)
For the higher order frame potential, a lower bound can be obtained as
$F^{(K)}=(1-q_{K})\left(\frac{N_{\rm s}^{2}N_{\rm b}}{N_{\rm s}^{2}N_{\rm b}^{2}-1}\right)^{n}-q_{K},$ (9)
where the factor ($K\geqslant 2$) is
$q_{K}=\left(1+\frac{2^{K}-1}{N_{\rm b}}\right)F_{\rm Haar}^{(K)}.$ (10)
Although Hamiltonian evolution, which is spatially local, differs from random Haar-sampled unitaries, it agrees well with the Haar unitary results when the number of measurements $n$ is small. As illustrated by the blue dotted lines in Fig. 3, the frame potential exhibits a rapid decay in this regime. This behavior is well captured by the HDT theory described in Eq. (7), which predicts an exponential decrease, as shown by the orange lines in Fig. 3. Notably, this agreement holds across different Hamiltonians, as demonstrated in Fig. 3(a)(b)(c).
Phenomenologically, when the frame potential approaches the same order as its minimal value, $F^{(1)}=(r+1)F_{\rm c}$, saturation starts to kick in. The number of measurements required for saturation is
$n_{\rm sat}\approx\frac{n_{\rm s}-{\rm log}_{2}r}{n_{\rm b}}\leavevmode\nobreak\ .$ (11)
After the saturation of the frame potential, small fluctuations remain. These are primarily due to finite-size effects and the limited evolution time in the Hamiltonian dynamics. A detailed explanation can be found in Appendix A.
Figure 4: Frame potential for $K$-designs corresponding to (a) $K=1$, (b) $K=3$, and (c) $K=10$. The black solid line represents the analytical approach. Blue dots indicate results from the non-integrable Hamiltonian ($J_{z}\neq 0$), while green dots correspond to the integrable Hamiltonian ($J_{z}=0$).
The dynamics of integrable and non-integrable systems are fundamentally different, and this distinction is also reflected in the behavior of the low-order frame potential. By tuning the Hamiltonian in Eq. (3), we can control its integrability. Specifically, setting $J_{z}=0$ renders the system integrable, whereas a nonzero $J_{z}$ leads to non-integrable dynamics.
In the regime of a small number of measurements $n$, the non-integrable system more closely follows the predictions of HDT, while the integrable system deviates more rapidly from the HDT curve, as shown in Fig. 4(a). As $n$ increases, both systems exhibit oscillatory behavior. However, as the order of the frame potential increases, the distinction between integrable and non-integrable systems becomes less pronounced, as illustrated in Fig. 4 (b) and (c).
For higher-order frame potentials, both systems saturate at similar values, and the oscillations gradually vanish as the design increases. This suggests that at high design, the distinction between integrable and non-integrable local Hamiltonians becomes negligible.
We find that both integrable and non-integrable local Hamiltonians remain far from the minimum value predicted by HDT theory, as shown in Fig. 4(c). This deviation is likely due to the inherent locality of the Hamiltonian. Although a chaotic Hamiltonian may exhibit ergodic behavior in local observables, it does not fully explore the entire Hilbert space. One fundamental limitation is energy conservation, which restricts dynamics to a finite energy shell. Moreover, due to locality constraints, the evolution cannot uniformly cover Hilbert space in the same way as a Haar random unitary. Furthermore, the Hamiltonian contains only a limited number of independent parameters, meaning that the matrix elements of the evolution operator $U=e^{-iH\Delta t}$ are not fully independent. In contrast, a Haar-random unitary has a number of independent parameters equal to the dimension of the Hilbert space minus one. Consequently, the dynamics generated by a local Hamiltonian cannot reach the level of randomness characteristic of a typical Haar-random unitary. These observations highlight the intrinsic limitations of using local Hamiltonians and restricted resources to generate high-order quantum randomness.
V Zeno effect
Excessively frequent measurements can freeze the quantum state due to the Zeno effect, thereby suppressing its inherent randomness. For a finite evolution time $T$, if the number of measurements $n$ is too large, the bath will remain almost fixed in its reset state $|\phi_{0}\rangle$. Consequently, the probability of obtaining any measurement outcome other than the reset state approaches zero. The system therefore remains in a pure state, with the frame potential approaching one.
The short-time dynamics can be analytically calculated in the limit of $\Delta t\rightarrow 0$, which allows us to investigate the asymptotic behavior of the frame potential.
Lemma 1
For a fixed total evolution time $T$, in the limit of $n\gg 1$, we obtain the following asymptotic lower bounds for the frame potential
$F^{(K)}\gtrsim e^{-O\left(\frac{T^{\alpha+1}}{n^{\alpha}}\right)},$ (12)
where $\alpha=1$ is a constant.
[…]
When $n$ is large, the frame potential approaches 1. To extract the exponential decay behavior, we plot the data on a logarithmic scale. The asymptotic behavior of the frame potential for different Hamiltonians is shown in Fig. 5. The numerical results are represented by the blue curves, while the exponential decay rate given in Eq. (12) is indicated by the green lines. As $n$ increases, the numerical results converge more closely to the predicted exponential behavior. In the case where $\left[H_{{\rm c}0},|\phi_{m}\rangle\langle\phi_{m}|\right]=0$, the decay exponent is $\alpha=3$, as illustrated in Fig. 5(a). Otherwise, when the commutator is nonzero, the exponent is $\alpha=1$, as shown in Fig. 5(b) and (c).
The starting point for the number of measurements required to observe the Zeno effect is approximately
$n_{\rm Zeno}$ $\approx$ $\sqrt[\alpha]{\frac{c_{H}T^{\alpha+1}}{n_{\rm s}\ln 2}},$ (13)
where $c_{H}$ is a constant arising from the Hamiltonian, and its value will be derived in Appendix B. The number of measurements should not exceed this value to preserve states with a good level of randomness.
The saturation of the frame potential requires that $n_{\rm sat}<n_{\rm Zeno}$. In other words, the total evolution time must satisfy
$T>\sqrt[\alpha+1]{\frac{(n_{\rm s}-{\rm log}_{2}r)^{\alpha}n_{\rm s}\ln 2}{n_{\rm b}^{\alpha}c_{H}}}\leavevmode\nobreak\ .$ (14)
To achieve a Haar-random ensemble, the total Hamiltonian evolution time should be greater than this threshold.
When the total evolution time $T$ is sufficiently short such that $n_{\rm sat}>n_{\rm Zeno}$, there is a competition between HDT and Zeno effect. In this regime, the frame potential does not reach its saturation value and remains significantly above it. The minimum frame potential is attained at around
$n_{\gamma}\approx T\sqrt[\alpha+1]{\frac{c_{H}}{n_{\rm b}\ln 2}}.$ (15)
This corresponds to the recommended number of measurements under a limited total evolution time, constrained by experimental factors like coherence time.
Figure 5: Frame potential as a function of the number of measurements for (a) $H_{\rm Ising}$, (b) $H_{YY}$, and (c) $H_{XXX}$. Blue lines represent numerical results, while green lines correspond to analytical predictions for Zeno effect with (a) $\alpha=3$, and (b)(c) $\alpha=1$. The total evolution time is $T=15$. The system consists of 5 qubits, with 3 bath qubit.
VI conclusion
We employ a protocol involving multiple Hamiltonian evolutions, measurements, and resets to access holographic deep thermalization. To quantify the degree of randomness generated, we use the frame potential. In the regime of few measurements, the frame potential decreases exponentially until it saturates at a minimum value, indicating maximal scrambling. In contrast, when measurements are too frequent, the frame potential rises exponentially because the quantum Zeno effect suppresses thermalization. We derive the asymptotic behavior of the frame potential analytically, which shows excellent agreement with our numerical results.
Acknowledgements.
The authors acknowledges discussions with Bingzhi Zhang. The project is supported by Office of Naval Research Grant No. N00014-23-1-2296 and DARPA HR0011-24-9-0362. QZ also acknowledges support from NSF (OMA-2326746, 2350153, CCF-2240641), AFOSR MURI FA9550-24-1-0349, DARPA HR0011-24-9-0362, Halliburton Company an unrestricted gift from Google.
Appendix A Revival dynamics
As the frame potential approaches its saturation value, small fluctuations appear, primarily arising from revivals in the Hamiltonian evolution of the finite-size system. In this regime, the energy spectrum becomes discrete with only a small number of energy levels, and the smoothing effect of thermalization diminishes.
To quantify these revivals, we use the fidelity between the evolved state $\psi_{t}$ and the initial state $\psi_{0}$, defined as
$\xi=\left|\langle\psi_{0}|\psi_{t}\rangle\right|^{2}.$ (16)
[…]
Figure 6: Fidelity of representative states during the time-dependent revival dynamics governed by $H_{\rm Ising}$. The system comprises 7 qubits and 1 bath qubit.
Appendix B Zeno effect
In the following, we derive Eq. (12). The Hamiltonian can be decomposed into three parts
$H$ $=H_{{\rm s}}+H_{{\rm b}}+H_{{\rm c}}$ (17)
$=H_{{\rm s}0}+H_{{\rm b}}+H_{{\rm c}0},$ (18)
where $H_{{\rm s}},H_{{\rm b}}$ and $H_{{\rm c}}$ corresponds to the system, bath and interaction Hamiltonian. In the second step, we introduced a modified system and interaction Hamiltonian
$H_{{\rm s}0}=H_{{\rm s}}+\langle\phi_{0}|H_{\rm c}|\phi_{0}\rangle,$ (19)
and
$H_{{\rm c}0}=H_{\rm c}-\langle\phi_{0}|H_{\rm c}|\phi_{0}\rangle\leavevmode\nobreak\ ,$ (20)
[…]
B.1 Low order Zeno effect
In the general case, the powers is $\alpha=1$. The measurement basis are changed under the operation $H_{{\rm c}0}$, Namely $\left[H_{{\rm c}0},|\phi_{m}\rangle\langle\phi_{m}|\right]\neq 0$. We can decompose the system wave function in the orthogonal basis
$|\psi_{t}\rangle=\sum_{l}a_{l}|\psi_{l}\rangle.$ (21)
We define the time scale
$\omega_{l^{\prime},m}=\sum_{l}a_{l}\langle\phi_{m}|\langle\psi_{l^{\prime}}|H|\psi_{l}\rangle|\phi_{0}\rangle.$ (22)
Although $\omega_{l^{\prime},m}$ depends on the initial system state $|\psi_{t}\rangle$, we can approximate it by replacing it with its average, i.e., $|\omega_{l^{\prime},m}|^{2}\rightarrow\overline{|\omega_{l^{\prime},m}|^{2}}$. In the deep thermalized regime, where $\Delta t$ is large, this average is effectively taken over random states. In contrast, when $\Delta t\rightarrow 0$, measurements occur so frequently that the bath remains effectively frozen in its initial state $|\phi_{0}\rangle$. In this limit, the system state $|\psi_{t}\rangle$ can be approximated as evolving under the unitary $e^{-iH_{\rm s0}t}$, and the average can be performed over pure time evolution.
In the limit of short-time evolution $\Delta t\rightarrow 0$, according to the Schrödinger equation, the probability of measuring the bath in the state $|\phi_{m}\rangle$ for $m\neq 0$ is
$P_{m}\approx\Delta t^{2}\sum_{l^{\prime}}\overline{|\omega_{l^{\prime},m}|^{2}}.$ (23)
The probability of the bath remain in state $|\phi_{0}\rangle$ is
$P_{0}=1-\sum_{m\neq 0}P_{m}\approx 1-\Delta t^{2}\sum_{l^{\prime}}\overline{|\omega_{l^{\prime},m}|^{2}}.$ (24)
For a rough lower bound, we retain only one term in the summation where all measurement results yield $|\phi_{0}\rangle$. The frame potential then satisfies
$F^{(K)}$ $>$ $P_{0}^{2n}$ (25)
$>$ ${\rm Exp}\left(-\frac{2T^{2}}{n}\sum_{l^{\prime}}\overline{|\omega_{l^{\prime},m}|^{2}}\right).$
In this case, the coefficient is $c_{H}=2\sum_{l^{\prime}}\overline{|\omega_{l^{\prime},m}|^{2}}$.
B.2 Hight order Zeno effect
In the special case, the powers is $\alpha=3$. The measurement basis $|\phi_{m}\rangle$} is invariant under the action of $H_{\rm c0}$—that is, when $\left[H_{\rm c0},|\phi_{0}\rangle\langle\phi_{0}|\right]=0$—we can derive a tighter lower bound on the frame potential.
To facilitate this derivation, we also decompose the bath Hamiltonian into a component that is diagonal in the measurement basis and a remainder term.
$H_{{\rm b}}=H_{{\rm bm}}+V,$ (26)
where $H_{{\rm bm}}$ commutes with the measurement.
In the limit of short time evolution $\Delta t\rightarrow 0$, the unitary could be written as [18]
$e^{-iH\Delta t}\approx e^{-iH_{{\rm c0}}\frac{\Delta t}{2}}e^{-iH_{{\rm s0}}\Delta t-iH_{{\rm b}}\Delta t}e^{-iH_{{\rm c0}}\frac{\Delta t}{2}}.$ (27)
[…]
$|\Psi\rangle$ $\approx$ $e^{-iH_{{\rm c0}}\frac{\Delta t}{2}}e^{-iH_{{\rm s0}}\Delta t-iH_{{\rm b}}\Delta t}e^{-iH_{{\rm c0}}\frac{\Delta t}{2}}|\psi_{t}\rangle|\phi_{0}\rangle$ (28)
$=$ $e^{-iH_{{\rm c0}}\frac{\Delta t}{2}}|\psi_{t+\Delta t}\rangle|\phi_{\Delta t}\rangle,$
where $|\psi\rangle$ represents the system wave function, while $|\phi\rangle$ denotes the bath wave function.
Expanding the bath wave function in the measurement basis $|\phi_{m}\rangle$, we have
$|\Psi\rangle\approx\sum_{m}e^{-i\langle\phi_{m}|H_{\rm c0}|\phi_{m}\rangle\frac{\Delta t}{2}}c_{m}|\psi_{t+\Delta t}\rangle|\phi_{m}\rangle.$ (29)
In the first-order time-dependent perturbation theory, we consider only the states directly coupled to the initial state. The corresponding amplitude is given by
$c_{m}$ $\approx$ $-\frac{1}{E_{0}-E_{m}}\langle\phi_{m}|V|\phi_{0}\rangle\left(e^{-i\left(E_{0}-E_{m}\right)\Delta t}-1\right),$
$P_{m}$ $\approx$ $4\left|\frac{\langle\phi_{m}|V|\phi_{0}\rangle}{E_{0}-E_{m}}\right|^{2}\sin^{2}\frac{\left(E_{0}-E_{m}\right)\Delta t}{2}.$ (30)
The frame potential can be simplified to
$F^{(k)}$ $>$ $\sum_{z,z^{\prime}}p_{z}p_{z^{\prime}}\left|\langle\psi_{z}|\psi_{T}\rangle\langle\psi_{T}|\psi_{z^{\prime}}\rangle\right|^{2K}$ (31)
$>$ $\left(\sum_{z}p_{z}\left|\langle\psi_{z}|\psi_{T}\rangle\right|^{2K}\right)^{2},$
where $|\psi_{T}\rangle$ is the system wave function, and the measurement results remain $|\phi_{0}\rangle$ throughout the entire process.
We can define the corrected factor
$P_{m}^{\prime}>P_{m}\cos^{2K}\frac{\mu_{m}\Delta t}{2},$ (32)
where $2\mu_{m}$ is the difference between the largest and smallest eigenvalues of $\langle\phi_{m}|H_{\rm c0}|\phi_{m}\rangle$.
The frame potential
$F^{(K)}$ $>$ $\left(\sum_{m}P_{m}^{\prime}\right)^{2n}$ (33)
$>$ $\left(1-\frac{1}{4}\sum_{m}\left|\langle\phi_{m}|V|\phi_{0}\rangle\right|^{2}K\mu_{m}^{2}\Delta t^{4}\right)^{2n}$
$>$ ${\rm Exp}\left(-\frac{KT^{4}}{2n^{3}}\sum_{m}\left|\langle\phi_{m}|V|\phi_{0}\rangle\right|^{2}\mu_{m}^{2}\right).$
In this case, the coefficient is $c_{H}=(K/2)\sum_{m}\left|\langle\phi_{m}|V|\phi_{0}\rangle\right|^{2}\mu_{m}^{2}$. Notice that $\mu_{m}=0$ if $\langle\phi_{m}|H_{\rm c0}|\phi_{m}\rangle=0$. Consequently, only a few terms contribute to the summation.